OpenStudy (anonymous):

Medal+Fan for explanation. Solve this equation. 1+sqrt3x+3=sqrt7x+2

OpenStudy (anonymous):

\[1+\sqrt{3x+3}=\sqrt{7x+2}\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

\[1+\sqrt{3x+3}=\sqrt{7x+2}\] squaring both sides,\[1+2\sqrt{3x+3}+(3x+3)=7x+2\]isloating the remaining root.\[2\sqrt{3x+3}=7x+2-(3x+3)-1\]

OpenStudy (anonymous):

can you continue from here? start from adding like terms on both sides.

OpenStudy (anonymous):

Okay, let me see, one second. Could you check what I do?

OpenStudy (anonymous):

sure sure...

OpenStudy (anonymous):

can you please post any work you have here. if you have any questions, then please ask.

OpenStudy (anonymous):

\[2\sqrt{3x+3}=7x+2-(3x+3)-1\]\[2\sqrt{3x+3}=7x+2-3x-3-1\]

OpenStudy (anonymous):

\[2\sqrt{3x+3}=4x-2\]

OpenStudy (anonymous):

Okay so then 2sqrt3x+3=4x+4?

OpenStudy (anonymous):

why +4?

OpenStudy (anonymous):

2+3-1?

OpenStudy (anonymous):

it is negative 3.

OpenStudy (anonymous):

I didn't distribute, my bad.

OpenStudy (anonymous):

Som after that we divide both sides by 2. \[\sqrt{3x+3}=2x-1\]

OpenStudy (anonymous):

\[3x+3=4x^2-4x+1\]

OpenStudy (anonymous):

\[0=4x^2-7x-2\]

OpenStudy (anonymous):

Subtract out the left side, I'm following you.

OpenStudy (anonymous):

lets test the discriminant. can you tell me what the dsicrimant is going to be?

OpenStudy (anonymous):

You mean the factor correct?

OpenStudy (anonymous):

no, I mean that you should (please)m find the discriminant for this. if the discriminant is a perfect square, then the polynomial is factorable.

OpenStudy (anonymous):

Well, it's not a perfect square.

OpenStudy (anonymous):

\[(-7)^2-4(1)(-2)~~\Rightarrow~~49+8~~\Rightarrow~~57\]

OpenStudy (anonymous):

you will have to use a quadratic formula thne.

OpenStudy (anonymous):

can you fidn what x is?

OpenStudy (anonymous):

by using the quadratic formula on 4x^2-7x-2

OpenStudy (anonymous):

ohh, I did it incorrectly.

OpenStudy (anonymous):

I mean my discriminant is, \[(-7)^2-4(-4)(2)~~\Rightarrow~~49+32~~\Rightarrow~~81\]

OpenStudy (anonymous):

it is a perfect square.

OpenStudy (anonymous):

THANK GOODNESS! That's what I got after going back over it!

OpenStudy (anonymous):

Okay, reminding the equation we have is,\[4x^2-7x+2=0\]

OpenStudy (anonymous):

can you factor it?

OpenStudy (anonymous):

One sec.

OpenStudy (anonymous):

sure:)

OpenStudy (anonymous):

Okay, this is quite confusing me. Can you assist?

OpenStudy (anonymous):

fill in blanks.\[(4x~~~~~~~~~~~~)(x~~~~~~~~~~~~)\]

OpenStudy (anonymous):

-7 +2? Doesn't it stay the same?

OpenStudy (anonymous):

whay -7 and +2?

OpenStudy (anonymous):

you need it to be, \[4x^2-7x+2=0\]

OpenStudy (anonymous):

I'm getting frustrated, factoring is the thing I'm quite worst at.

OpenStudy (anonymous):

\[4x^2-7x+2=0\]\[(4x+1)(x-2)=0\]

OpenStudy (anonymous):

can you tell me what x equals?

OpenStudy (anonymous):

Yeah, I can do that.

OpenStudy (anonymous):

4x+1=0 which means 4x=-1 OR x=-1/4 AND x-2=0 which means x=2

OpenStudy (anonymous):

yes, x=-1/4 is one solution, and x=2.

OpenStudy (anonymous):

Gr. Okay.

OpenStudy (anonymous):

if you plug in -1/4 into square roots, do you get any negative numbers in the roots, or not?

OpenStudy (anonymous):

No. Because a negative times a negative is never a negative right?

OpenStudy (anonymous):

you need to make sure that when x=-/14 the square roots of your original equation are not negative

OpenStudy (anonymous):

Oh, I see where you are. I'm sorry I'm so slow today. MAth is killing me. One moment.

OpenStudy (anonymous):

take your time.

OpenStudy (anonymous):

Uh, yeah they're negative. Duh.

OpenStudy (anonymous):

The first isn't.

OpenStudy (anonymous):

Oh wait. Neither is the second false alarm.

OpenStudy (anonymous):

yes, none of the roots are negative when x=-1/4.

OpenStudy (anonymous):

But it was important to check.... b/c we might have gootn some imaginary numbers.

OpenStudy (anonymous):

So what does that mean?

OpenStudy (anonymous):

Ohh, I see I see.

OpenStudy (anonymous):

that x=-1/4 and x=2 are both fine solutions.

OpenStudy (anonymous):

Thank you.. Do you mind one more?

OpenStudy (anonymous):

okay, but I might have to leave at any second.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Divide and Simplify if possible sqrt of 250x^16/sqrt2x

OpenStudy (anonymous):

\[\Large \frac{\sqrt{250x^{16}}}{\sqrt{2x}}\]yes?

OpenStudy (anonymous):

Exactly.

OpenStudy (anonymous):

first, multiply times sqrtP2x{ on top and vbottom, to rationalize it.

OpenStudy (anonymous):

sqrt{2x}

OpenStudy (anonymous):

can you do this?

OpenStudy (anonymous):

Nope. Listen if you're in a rush I can get someone else in /.\ don't wanna be a bother

OpenStudy (anonymous):

I got to go, I am in arush... and that is nice of you:) bye, and hope to see you some time later.